OpenStudy (ihelpyoulearn):

Who here is willing to help me out with some inequalities questions? I will, like always, fan and medal everyone who helps :)

OpenStudy (mortonsalt):

Which questions, rather, is the better question haha

OpenStudy (ihelpyoulearn):

Well 1. Which number is a solution of the inequality? 10.6<b -18 -9 7 14

OpenStudy (ihelpyoulearn):

2. Which number is a solution of the inequality? g+3>6 3 4 -3 -4

OpenStudy (hwyl):

WHAT YOU NEED IS A LESSON HOW TO UNDERSTAND AND WORK ON THOSE PROBLEMS, NOT ANSWERS TO YOUR PROBLEMS

OpenStudy (ihelpyoulearn):

I know what OpenStudy is for!

OpenStudy (ihelpyoulearn):

Closing question. I guess I can help others all day but never get help back :-/ Oh well.

OpenStudy (hwyl):

GIVING ANSWERS IS NOT HELPING

OpenStudy (ihelpyoulearn):

@hwyl I didn't ask for answers! I asked for 'help' on inequality questions. So not sure where you got I'm asking for full out answers! I'm not going to argue like a little kid over something stupid. Next time, if you aren't going to help don't reply :) Easy!

OpenStudy (mortonsalt):

I'm still willing to help if that counts for anything. Before you'll be able to solve inequalities, you have to recognise the different symbols involved. Familiarise yourself with the less than, greater than, and equal signs. Next, when you solve an inequality, your goal is to isolate the variable to one side of the equation. For your first question, b is already isolated to one side. This part now is a bit more of common sense. According to the equation "b" has to be greater than 10.6. Which of the four numbers is greater than 10.6?

OpenStudy (mortonsalt):

For question 2, just treat the inequality like you would a regular equation. With that, we can say: \[g+3 > 6 \] \[g>6-3\] \[g>3\] To check your answer, just plug values that are GREATER than 3 into the original inequality. (Remember: GREATER, not equal to. Be careful with the signs because they could easily be overlook.)

OpenStudy (mortonsalt):

Generally, these are the basics to solving inequalities. The "position" at which the symbol is facing would not change unless: 1. You are dividing by a negative number on both sides. 2. You're intentionally swapping the right and left hand sides. For 1, here's an example: \[-2x>3\] When we divide both sides by -2, you have to switch the signs. \[\frac{-2x}{-2} > \frac{3}{-2}\] \[x<\frac{-3}{2}\] To check your answer, try plugging a number that is less than -3/2 (e.g. -4/2) to the original inequality. If your inequality is false after the checking, then you might have missed a step. If it is true, you're good to go!

OpenStudy (ihelpyoulearn):

@mortonsalt Thank You so much! This really helped me understand :) I appreciate it!